Consider the following monopolist who has 10 customers with the following willingness to pay (all willingness to pay in in $ per unit of the good): 10, 20, 30, 40, 50, 60, 70, 80, 90, 100. The marginal cost of the monopolist is $25/ unit of the good. a) If the monopolist can perfectly price discriminate, then find out how many units will the monopolist sell. Also calculate the profit of the monopolist. [3+4]
Answer #1

If monopolist will Perfectly price discriminate, he will charge the maximum willingness to pay by all customers until the willingness to pay is more than the marginal cost

Hence all customers with more than $25 willingness to pay will be served. Hence 8 units will be sold.

Hence 8 units is answer

Profit from $100 customer=100-25=$75

Profit from $90 customer=90-25=$65

Profit from $80 customer=80-25=$55

Profit from $70 customer=70-25=$45

Profit from $60 customer=60-25=$35

Profit from $50 customer=50-25=$25

Profit from $40 customer=40-25=$15

Profit from $30 customer=30-25=$5

Total profit=75+65+55+45+35+25+15+5=$320